OpenStudy (ttop0816):

Which of the following equations are equivalent to -2m - 5m - 8 = 3 + (-7) + m? -15m = -4m -7m - 8 = m - 4 -3m - 8 = 4 - m m - 4 = -7m - 8 -8 - 7m = -4 + m -8 - 3m = 4 - m

OpenStudy (ttop0816):

it can have multiple answers!

OpenStudy (anonymous):

-7m - 8 = 4 - m m - 4 = -7m-8 -8 - 7m = -4 + m

OpenStudy (ttop0816):

@badugu2008 thank you! do you mind helping me out with another question?

OpenStudy (ttop0816):

i know its really simple but i just got a little bit confused! Solve -3 = 1/2x + 1/2x + x. -5 1/2 -1 3/4 -1 1/2 Solve 7 + x - 15 = -2 2/3. 5 1/3 6 1/3 10 2/3 20 2/3

OpenStudy (anonymous):

-1 1/2 5 1/3

OpenStudy (ttop0816):

thank you so much!